OpenStudy (anonymous):

A square has an area of 16 cm². Calculate the length of its diagonal to the nearest 10th of a centimeter. A. 6.4 cm B. 5.7 cm C. 8 cm D. 4 centimeters E. 64 cm

OpenStudy (campbell_st):

draw a diagram  area 16cm^2 gives side length 4cm use pythagoras' theorem 4^2 + 4^2 = x^2 x^2 = 32 \[x = \sqrt{32} = 4\sqrt{2}\] you'll need to evaluate to find x